OpenStudy (sophadof):

What is the sum of 2/x+4/x+1=9

OpenStudy (dayakar):

\[\frac{ 2 }{ x }+ \frac{ 4 }{ x }+1=9\] is it what u asked?

OpenStudy (sophadof):

\[\frac{ 2 }{ x }+\frac{ 4 }{ x+1 }=9\]

OpenStudy (sophadof):

@dayakar sorry about that!

OpenStudy (dayakar):

ok

OpenStudy (dayakar):



OpenStudy (dayakar):

can u do it remaining part

OpenStudy (sophadof):

sorry.. im trying it!!(: @dayakar

OpenStudy (dayakar):

do u understand above step

OpenStudy (sophadof):

yes! @dayakar

OpenStudy (sophadof):

but see I thought it was 1/3 but it's not Here are the choices a.-1/3 b.0 c.1/3 d.1 @dayakar

OpenStudy (dayakar):

do u want answer complete explanation

OpenStudy (sophadof):

I just need the answer(: unless you want to give the explanation but I really don't need it.. @dayakar

OpenStudy (dayakar):

substitute x= 1/3 in original problem then u will get both sides equal number

OpenStudy (sophadof):

Im trying right now @dayakar

OpenStudy (dayakar):

show your work

OpenStudy (sophadof):

it should be 1/3 why is the test saying im wrong though... @dayakar

OpenStudy (sophadof):

I even checked it

OpenStudy (dayakar):

i am doing step by step 2/x = 2/(1/3) = 2* 3/1 = 6 is it clear

OpenStudy (dayakar):

x+1 = 1/3 +1 = 1/3 +3/3 [ 1= 3/3] = (1+3)/3 = 4/3

OpenStudy (dayakar):

4/(x+1) = 4/(4/3) = 4*3/4 = 3 [ 4 cancelled]

OpenStudy (dayakar):

2/x + 4/(x+1) = 9 {given problem] 6 + 3 = 9 [ left hand side part we did above] 9 =9 [ true ] our answer is 1/3